Paul is twice as old as Becca. Five years ago the sum of their ages was 44. How old are they now?
How do I set this up to get my answer?

Answer by ewatrrr(24785)

x Becca's age (Paul 2x)
5 years ago...
(2x-5) + (x-5) = 44
x = 54/3 = 18, Becca's age. Paul 36
And..checking
13 + 31 = 44